# GPT-4V with Emotion: A Zero-shot Benchmark for Generalized Emotion Recognition

### 03 Feb 2024 v2

@@ 6 Conclusion
This paper conducts a comprehensive evaluation of GPT-4V in GER tasks. It is worth noting that GPT-4V exhibits strong visual understanding capabilities, even surpassing supervised systems in visual sentiment analysis. However, it performs poorly in micro-expression recognition which requires specialized knowledge. This paper also showcases GPT-4V’s temporal modeling and multimodal fusion capabilities, along with its robustness to color space changes. Furthermore, we evaluate its prediction consistency and security check stability, and visualize error cases to reveal its limitations in emotion understanding. This paper can serve as a zero-shot benchmark and provide guidance for subsequent research on emotion recognition and multimodal LLMs. In the future, we will expand the evaluation scope by including more emotion-related tasks and datasets.
@@ Acknowledgements
This work is supported by the National Natural Science Foundation of China (NSFC) (No.62201572, No.61831022, No.62276259, No.U21B2010), Beijing Municipal Science & Technology Commission, Administrative Commission of Zhongguancun Science Park No.Z211100004821013, Open Research Projects of Zhejiang Lab (NO. 2021KH0AB06), and CCF-Baidu Open Fund (No.OF2022025).

### 18 Mar 2024 v3

@@ 7 Conclusion
This paper conducts a comprehensive evaluation of GPT-4V in GER tasks. Our evaluation results demonstrate that GPT-4V exhibits strong visual understanding capabilities and achieves promising results in general-purpose emotion recognition. However, it performs poorly in micro-expression recognition which requires specialized knowledge. This paper also showcases GPT-4V’s temporal modeling and multimodal fusion capabilities, along with its robustness to color space and prompt template changes. Furthermore, we evaluate its prediction consistency and security check stability and visualize error cases to reveal its limitations in emotion understanding. This paper can serve as a zero-shot benchmark and provide guidance for subsequent research on emotion recognition and MLLMs.
@@ Acknowledgements
This work is supported by the National Natural Science Foundation of China (NSFC) (No.62201572, No.61831022, No.62276259, No.62322120, No.U21B2010) and Beijing Municipal Science & Technology Commission, Administrative Commission of Zhongguancun Science Park (No.Z211100004821013).